OpenStudy (anonymous):

A battery is 10% charged and is charging at a constant rate. The graph shows the battery’s charge over time. What is the slope of the line that represents the situation? Enter the slope as a fraction. First quadrant grid, x-axis numbered 0 to 130 and labeled Minutes, y-axis numbered 0 to 100 and labeled Percent charged. Line passes through labeled points at 40, 40) and (120, 100).

OpenStudy (alexandervonhumboldt2):

slope is (40-100)/(40-120)

OpenStudy (alexandervonhumboldt2):

-60/-80=6/8=????????

OpenStudy (alexandervonhumboldt2):

3/4 slope is 3/4

OpenStudy (anonymous):

thank you so much do you mind helping me with just a few more

OpenStudy (alexandervonhumboldt2):

i can help can you plz medl me?

OpenStudy (anonymous):

Manuel builds and sells wooden crates. He made a table to show how much his crates cost. Which equation represents the situation? Manuel’s Crates Crates 2 4 6 8 Cost 25 45 65 85 A. y – 2 = −10(x – 25) B. y – 4 = 10(x – 45) C. y – 65 = −10(x – 6) D. y – 85 = 10(x – 8)

OpenStudy (alexandervonhumboldt2):

so slope intersept is y=10x+5